### v2

# The NordDRG AI Benchmark for Large Language Models

The NordDRG AI Benchmark for Large Language Models
Tapio Pitkäranta
Department of Computer Science and Engineering
Aalto University, Finland
tapio.pitkaranta@iki.fi
Large-language models (LLMs) are already being piloted for clinical coding and decision support, yet—until now—no open benchmark has targeted the hospital-funding layer where Diagnosis-Related Groups (DRG) determine reimbursement across many countries. We release NordDRG-AI-Benchmark, the first public test-bed that captures a complete DRG rule set and probes an LLM’s ability to reason over multilingual diagnosis, procedure, and tariff logic.
[…]
All artefacts are released at GitHub  [^1] https://github.com/longshoreforrest/norddrg-ai-benchmark.
A baseline demonstration shows that five state of the art LLMs diverge markedly on the nine automatically verifiable tasks: o3 (OpenAI) scores 9 / 9, GPT-4o and o4-mini-high 7 / 9, while Gemini 2.5 Pro and Gemini 2.5 Flash solve only 5 / 9 and 3 / 9 respectively. These results confirm that NordDRG-AI-Benchmark surfaces domain-specific strengths and weaknesses that remain hidden on generic LLM leaderboards, and thus provides a reproducible yard-stick for future research on trustworthy automation in hospital funding.
Keywords Artificial Intelligence AI, Large Language Models LLM, Diagnosis Related Groups DRG, NordDRG, CaseMix systems, Clinical classifications
1 Introduction
1.1 Problem Identification
Large-scale generative models have progressed from research prototypes to production-grade services in under five years, bringing natural-language interfaces to domains once considered too specialised for automated reasoning. Health-care finance is a notable example: every inpatient episode in most OECD countries is mapped to a Diagnosis-Related Group (DRG) code that determines reimbursement, yet the rule sets governing these allocations remain opaque to non-experts and are revised annually through lengthy committee processes. Despite a surge in medical-LLM research, no open benchmark currently targets this socio-technical bottleneck—creating a blind spot for both the AI community and hospital administrators who need trustworthy decision support.
1.2 Contributions
We address this shortfall by releasing NordDRG-AI-Benchmark, the first public benchmark that unites clinical coding, multilingual reasoning and health-care finance. Specifically, we contribute (i) a machine-readable dump of the NordDRG definition tables, (ii) governance manuals that capture real-world change-management workflows, and (iii) 14 task-oriented prompts spanning code lookup, cross-table inference and qualitative audit scenarios. All artefacts are released to catalyse research on transparent, regulator-compliant language models for hospital funding.
1.3 Research methodology.
The work is grounded in the Design Science Research (DSR) paradigm, which couples rigorous inquiry with artefact construction to solve practical problems while contributing to theory. Following the canonical six-step model of Peffers et al. [Peffers et al., 2020], we (1) identified the absence of a CaseMix benchmark as a critical application–evaluation gap, (2) specified objectives for a reproducible LLM test-bed, (3) designed and built the NordDRG artefact suite, (4) demonstrated its utility through 14 task scenarios, (5) evaluated baseline performance with selected LLMs, and (6) communicated the results via this open-access release. Two rapid build–evaluate cycles with domain experts ensured that the artefacts remain aligned with real-world reimbursement workflows.
Author note: LLM use  [^2] During manuscript preparation the authors made limited use of OpenAI services (GPT-4o, o3, and o4-mini) for grammar, stylistic refinement, and reference-list formatting. All prompts, intermediate outputs, and final wording were reviewed and verified by the authors, who accept full responsibility for the content.
1.4 Research gap
Recent studies have begun to probe whether large-language models can predict U.S. MS-DRGs [Wang et al., 2023, Wang et al., 2025] or internalise isolated grouper rules [Hajialigol et al., 2023, He et al., 2022], yet the evidence base remains fragmented. All published evaluations (i) focus on a single national system, (ii) collapse the multi-table CaseMix logic into a flat classification task, and (iii) rely on private datasets that cannot be redistributed. Consequently, the community still lacks a public, multilingual benchmark that captures the complete rule graph of a CaseMix system such as NordDRG—including cross-table joins, governance documents, and annual change-log workflows. Without such a yard-stick, researchers cannot (a) establish reproducible baselines for LLM reasoning over reimbursement logic, (b) compare design patterns (fine-tuning, retrieval, reinforcement learning, chain-of-thought) under identical conditions, or (c) trace progress across model generations and academic groups.
These shortcomings crystallise into two open challenges:
  1. 1.
    Scope Gap — No prior work covers the full, multilingual CaseMix rule set used in Nordic hospital funding; existing DRG experiments are confined to reduced or proprietary subsets.
  2. 2.
    Benchmark Gap — The absence of an open, standardised CaseMix benchmark precludes rigorous, head-to-head comparison of alternative LLM architectures and hinders cumulative research on trustworthy automation in hospital finance.
1.5 Research objectives
Guided by the two gaps identified above, this study pursues three concrete objectives:
  1. 1.
    O1 — Artefact construction.  Design and release an open, machine-readable benchmark (NordDRG-AI-Benchmark) that captures the complete NordDRG rule graph—including multilingual group names, property tables, and governance manuals—and packages it in a modular format for easy reuse.
  2. 2.
    O2 — Task operationalisation.  Instantiate a balanced test-suite of 14 DRG CaseMix prompts (9 quantitative, 5 qualitative) that reflect real workflows of clinicians, coders, and policy analysts, together with a gold answer key for automatic or rubric-based scoring.
  3. 3.
    O3 — Baseline evaluation.  Provide first-cut results for five state-of-the-art LLM endpoints, thereby establishing a reproducible yardstick for future optimisation studies in hospital-funding contexts.
2 Background and Related Work
2.1 NordDRG CaseMix System
The Nordic Diagnosis-Related Groups (NordDRG) classification is the shared CaseMix framework used throughout the Nordic countries to cluster inpatient episodes with comparable resource consumption [NordicCasemixCentre, 2025]. By assigning each hospital stay to a single DRG, the system underpins transparent benchmarking of costs, outcomes, and efficiency across institutions, regions, and years.
NordDRG is distributed as a set of definition tables—about twenty inter-linked sheets released annually in Excel, CSV, and JSON formats—that encode the logic connecting diagnosis and procedure codes, age/sex splits, and national activation flags to individual DRG codes. These tables constitute the authoritative rule base for coding software, technical-change management, and policy analysis.
Because each rule specifies explicit inclusion criteria, identical cases are grouped consistently regardless of hospital or country. At the same time, each nation can activate, deactivate, or re-weight DRGs locally, allowing the common core to accommodate diverse clinical practices and tariff structures. This blend of rule-level precision and country-level flexibility makes NordDRG a durable instrument for hospital finance, clinical research, and health-system governance in the Nordic region.
2.2 Large Language Models (LLMs)
The Transformer architecture of Vaswani et al. [Vaswani et al., 2017] triggered the modern surge in natural-language processing. Early encoder models such as BERT [Devlin et al., 2018] were soon joined by large autoregressive generators including GPT-3 [Brown et al., 2020] and GPT-4, each scaling to hundreds of billions of parameters and exhibiting emergent abilities in text generation, retrieval, and reasoning. Parallel research explored alternative training paradigms (T5’s text-to-text framework [Raffel et al., 2020]), compute–data trade-offs (Chinchilla [Hoffmann et al., 2022]), and extreme scale (PaLM’s 540-billion-parameter Pathways model [Chowdhery et al., 2022]). In tandem, open-weight families such as OPT [Zhang et al., 2022] and Llama-2 [Touvron et al., 2023] widened community access to state-of-the-art capabilities. Recent reports even suggest that certain LLMs have passed variants of the Turing test [Jones and Bergen, 2025], underscoring their capacity for human-like communication. Comprehensive surveys chart this rapid evolution and situate LLMs within the broader arc of generative AI—from GANs to ChatGPT—highlighting key architectural and application trends [Naveed et al., 2023, Cao et al., 2023].
Since late 2022, chat-oriented front-ends—exemplified by ChatGPT—have carried LLMs from research labs into everyday workflows. The conversational interface abstracts away programming and domain expertise, turning LLMs into a general-purpose layer through which users in healthcare, law, education, and industry can delegate complex cognitive tasks.
As capabilities expand, the central question moves from can we do it? to can we do it safely?. Aligning model behaviour with human values and organisational goals has become a pressing agenda [Christian, 2020]. Achieving this alignment demands a multidisciplinary approach that blends data science with ethics, governance, and leadership [Wilson, 1999], ensuring that powerful generative models are deployed responsibly and transparently.
2.3 LLM Benchmarks and Leaderboards
Figure 1: LLM Leaderboard illustration: Vellum leaderboard
https://www.vellum.ai/llm-leaderboard
Benchmark suites and public leaderboards now play a decisive role in charting progress across the rapidly evolving LLM landscape. Early task collections such as GLUE [Wang et al., 2018] and SuperGLUE [Wang et al., 2019] have given way to multi-domain, multi-skill batteries—e.g., BIG-bench, [bench Collaboration, 2022], MMLU [Hendrycks et al., 2021], and the BabyLM Challenge [Warstadt et al., 2023]—that probe factual knowledge, mathematical reasoning, multilingual transfer, and instruction following. Complementary meta-evaluations such as HELM [Liang et al., 2022] standardise experimental settings and report robustness, bias, and efficiency metrics side by side.
Results from these test suites are aggregated on widely watched leaderboards. The HuggingFace Open LLM Leaderboard publishes automatic scores (e.g., MMLU, ARC, HellaSwag) for hundreds of open-weight models, while the LMSys Chatbot Arena [^3] https://chat.lmsys.org/ crowdsources pairwise human preferences to rank dialogue quality. Other platforms—including Vellum-AI [^4] https://www.vellum.ai/llm-leaderboard and MT-Bench [Zheng et al., 2023]—focus on application-specific criteria such as tool use or code generation. Figure 1 shows a representative snapshot of one such leaderboard, underscoring how transparent, regularly updated scorecards have become the de-facto currency for comparing models, spurring innovation, and diagnosing remaining failure modes.
2.4 Related Work: DRGs and LLMs
Large-language models (LLMs) are increasingly explored as end-to-end engines for hospital-reimbursement workflows. [Wang et al., 2023] fine-tune LLaMA on 236 k multimodal MIMIC-IV episodes, achieving 52 % top-1 accuracy over 771 MS-DRGs—a strong indication that the model captures the CC/MCC rules traditionally hard-coded in grouper software. Complementing this, [Kwan, 2024] show that a tool-augmented GPT-4 rivals human coders in ICD assignment, implying that careful retrieval and validation can offset the need for proprietary data. Zero-shot modelling is pushed further by [Renc et al., 2024], whose transformer infers DRG-like resource strata directly from longitudinal health trajectories, hinting at reimbursement systems emergent from latent patient-state geometry.
Beyond accuracy metrics, recent studies ask whether LLMs truly understand DRG logic. [Wang et al., 2025] introduce DRG-Sapphire, using reinforcement learning with rule-aware rewards and generating physician-validated rationales that explicitly cite grouping criteria. Interpretability is also central to [Hajialigol et al., 2023], whose multi-task transformer highlights token-level evidence so auditors can verify alignment with official rules. Structured knowledge helps as well: [He et al., 2022] inject a clinical knowledge graph into BERT and observe the largest gains on DRGs whose definitions hinge on multi-entity interactions, suggesting genuine rule-level reasoning. Finally, generic LLMs remain competitive; [Boyle et al., 2023] repurpose them for zero-/few-shot ICD coding by exploiting the ICD hierarchy, while [Boukhers et al., 2024] couple LLaMA representations with a lightweight MultiResCNN, reporting F1 gains across the full label set.
2.5 Our previous work
Our preliminary results appeared at [Pitkäranta, 2024b, Pitkäranta, 2024a] . The current article substantially extends that work by: (i) releasing a fully curated benchmark with 14 task-oriented prompts, (ii) providing open access github repository with all artefacts and, (iii) demonstrating baseline performance of five recent LLMs on the nine automatically verifiable tasks
3 Design & Development: NordDRG Benchmark
NordDRG-AI-Benchmark translates the vision set out in the Introduction into a fully modular test-bed. It packages the entire Nordic Diagnosis-Related Groups rule base—twenty inter-linked definition tables, annual governance manuals, and curated table excerpts—into machine-readable artefacts that can be loaded singly or in combination. A companion prompt pack of fourteen CaseMix tasks then probes three distinct facets of LLM capability:
  1. 1.
    Relational reasoning over complex DRG logic,
  2. 2.
    Multilingual understanding of clinical terminology, and
  3. 3.
    Procedural comprehension of the change-management workflows that drive yearly tariff updates across six NordDRG user countries.
By unifying these elements in a reproducible, open-source framework, the benchmark enables head-to-head evaluation of prompt engineering, retrieval augmentation, and fine-tuning strategies—offering, for the first time, a domain-specific yardstick for trustworthy automation in hospital funding.
All resources are released under at GitHub  [^5] https://github.com/longshoreforrest/norddrg-ai-benchmark and are grouped into four artefact classes:
  • A1 — Full definition corpus (.xlsx/.csv): roughly twenty relational sheets that encode the complete NordDRG grouper logic—diagnoses, NCSP procedures, age/sex splits, country-activation flags, surgery properties and grouping features.
  • A2 — Curated table excerpts (FI / FI–EN) (.xlsx/.csv): a Finnish subset released in two flavours—(i) Finnish-only labels and (ii) parallel Finnish–English labels—so researchers can probe LLM performance on minority-language data and bilingual transfer without loading the entire corpus.
  • A3 — Expert documentation (.pdf): governance manuals, change-log templates, and column-by-column guides that explain how technical updates are proposed, audited, and rolled into the annual NordDRG release.
  • A4 — CaseMix prompt pack (.xlsx): fourteen evaluation prompts aligned with Tables 34— nine quantitative tasks (IDs 1–9) with deterministic answer keys and five qualitative audits (IDs 10–14) that require free-text justification. Each prompt carries an Easy, Medium, or Hard tag and is paired with a gold-standard answer sheet.
Because every artefact is self-contained, users can mix and match them into four canonical “dataset configurations” (Table 2). This modular design lets researchers start with the lightweight Finnish-only excerpt for rapid prototyping, then scale up to the full tables plus governance PDFs to stress-test chain-of-thought reasoning and cross-document grounding.
Datasets
Table 1 gives a compact overview of the public resources that make up the benchmark. The paragraphs that follow unpack each entry—explaining how the definition tables, governance PDFs, curated Finnish / Finnish-English excerpts, and auxiliary prompts are constructed and how they can be combined in downstream experiments.
Table 1: Publicly released NordDRG benchmark resources.
Dataset Format Description
Definition Tables XLSX NordDRG Definition Tables
PDF Instructions PDF NordDRG PDF instructions
Definition Tables: Part of the tables FI XLSX NordDRG Definition Tables—Finnish DRG list with Finnish descriptions only
Definition Tables: Part of the tables FI–EN XLSX NordDRG Definition Tables—Finnish DRG list with Finnish and English descriptions of groups
Additional Prompts TXT NordDRG additional prompts
Evaluation scenarios.
The prompt pack crystallises 14 real-world CaseMix tasks that clinicians, health-economists, and coders routinely perform. Typical examples include
  • identifying DRGs for a specific procedure (e.g., appendectomy),
  • tracing diagnostic linkages for a rule such as 14X03,
  • checking whether drg_logic surgery flags align with proc_feat.
Nine grouping feature tasks (IDs 1–9) are quantitative: they admit deterministic answer keys and therefore support exact-match scoring. The remaining five tasks (IDs 10–14) are qualitative: they require free-text justifications or change-request drafts and are graded with a rubric-based human review.
Taken together, the fourteen prompts constitute a self-contained test suite for head-to-head comparisons of prompt engineering, retrieval augmentation, and fine-tuning strategies. By releasing this benchmark, we provide the first public, reproducible yardstick for LLMs in hospital-funding contexts, thereby closing partially the application and evaluation gaps identified earlier and laying the groundwork for transparent, regulator-compliant clinical language models.
3.1 Artefact: Definition Tables
[…]
  • drg_logic — rules that map diagnosis, procedure, age limits, CC/MCC status, and surgery flags to a unique DRG code; ordered by the ORD execution column.
  • dg_feat / proc_feat — full ICD-10 and NCSP diagnosis and procedure code catalogues annotated with properties referenced by drg_logic (e.g. DGCAT, COMPL, OR, PROCPRO).
  • Lookup sheets such as compl_excl, procprop_name, and dgprop_name, which define complication categories, OR-class mappings, and diagnosis property labels.
  • age_split and country_flag tables that hold paediatric/geriatric break-points and activation flags for each national version.
Each worksheet exposes stable identifier columns (e.g., DRG, ICD, NCSP, ORD), allowing the tables to be joined loss-lessly and extended by appending new rows in future releases. A representative excerpt is shown in Figure 2.
Figure 2: Materials used: NordDRG definition tables https://nordcase.org/forum/
3.2 Artefact: Expert documentation
Artefact 3.2 bundles the procedural knowledge that turns the raw tables into a living, governable system. Two official handbooks from the Nordic Casemix Centre—supplied as searchable PDFs—and the associated Excel template are included:
  • How to read the NordDRG definition tables A walkthrough of the Excel workbook that constitutes the grouper logic. The manual explains every high-impact sheet—including drg_logic, dg_feat, proc_feat and compl_excl—and shows how to (i) locate all rules that drive a given DRG, (ii) identify which diagnosis or procedure properties generate CC/MCC splits, and (iii) trace OR-flags, MDCs, age limits, and other row-level conditions.
  • How to write technical changes for NordDRG The canonical SOP for annual updates. It defines the “IN/OUT” convention, file-naming rules for TC_TEMPLATE_YYYY-MM-DD.xltx, and per-table editing guidelines—for example, duplicating a drg_logic row when altering ORD or ID, propagating new COMPL values consistently to dg_feat, compl_name, and compl_excl, and adding English code texts in the comments field for new ICD/NCSP codes.
Together, these documents specify (i) how existing grouping rules must be interpreted and (ii) how proposed modifications are documented, validated and merged. Including them in the benchmark ensures that LLM evaluations test not only code look-ups but also compliance with NordDRG’s governance process. A sample page is shown in Figure 3.
Figure 3: Materials used: NordDRG definition table expert documentation https://nordcase.org/forum/
3.3 Artefact: Definition Tables: Part of the tables FI
Artefact 3.3 distils the full workbook down to a single, lightweight file. The sheet contains only those DRG codes that are activated in the Finnish version and provides their official Finnish descriptions (DRG_TEXT_FI). Auxiliary columns preserve the DRG code, MDC and version stamp so that rows can still be cross-referenced against the master tables.
This compact subset supports two complementary use cases: first, it enables minority-language evaluation by allowing researchers to probe how well an LLM grasps Finnish DRG terminology— without loading the full 20-sheet corpus—and, second, it facilitates rapid prototyping, providing an ultra-light file for few-shot prompting, in-context search, or on-device experiments where the complete tables would exceed context limits.
3.4 Artefact: Definition Tables: Part of the tables FI-EN
Artefact 3.4 mirrors the row set of the Finnish-only file but appends the official English group descriptions (DRG_TEXT_EN). The bilingual layout lets LLMs align Finnish terminology with its English counterpart, enabling cross-lingual retrieval, code-switching prompts, and translation-quality checks within a single, lightweight worksheet. Such functionality is essential for multilingual healthcare settings, where clinicians, coders, and decision-support tools must seamlessly toggle between local and international nomenclatures
3.5 Artefact: Dataset configurations
Artefact 3.6 defines four ready-made modular configurations of the benchmark. Each primary artefact—definition tables, governance PDFs, and curated table excerpts—can be loaded on its own or combined, depending on the experiment. Table 2 lists four ready-made “dataset configurations” that map onto typical research needs.
The first pair, Definition Tables and Definition Tables + PDF Instructions, delivers the full-fidelity reference material used by hospital coders and is best suited for testing an LLM’s ability to reproduce the official NordDRG grouper logic or to justify decisions with citations from the manuals.
The second pair contains only a slice of the master tables: Finnish-only rows (FI) or bilingual Finnish–English rows (FI–EN). These lightweight subsets enable rapid prototyping and few-shot prompting in resource-constrained settings while still preserving clinically meaningful group descriptions.
Researchers can therefore start with the compact excerpts to establish baselines, then scale up to the full tables plus manuals to stress-test chain-of-thought reasoning, multilingual transfer, and cross-document grounding. Collectively the four bundles offer a flexible design space for analysing how data volume, document modality, and language coverage affect LLMs capability to comprehend NordDRG classification.
Table 2: Configuration types for the NordDRG benchmark datasets.
Dataset Configurations Type Description
Definition Tables Tables NordDRG Definition Tables
Definition Tables + PDF Instructions Tables+Instructions Definition tables and PDF descriptions
Definition Tables: Part of the tables FI PortionOfTables NordDRG Definition Tables: Finnish DRG list with Finnish descriptions only
Definition Tables: Part of the tables FI–EN PortionOfTables NordDRG Definition Tables: Finnish DRG list with Finnish and English descriptions of groups
3.6 Artefact: Additional Prompts
Artefact 3.6 supplies a single, illustrative prompt authored by NordDRG specialists (see Figure 4). The example walks an LLM through a four-step template—locate the relevant rows in drg_name, pull the multilingual descriptors, check country activation, and draft an English rationale—thereby showing how raw tables can be transformed into a reproducible reasoning task.
We ship only this seed prompt; it is intended as a blueprint that other researchers can copy, adapt, and extend to create their own task variants.
Figure 4: Materials used: NordDRG definition table additional prompts
3.7 Artefact: NordDRG CaseMix questions
Artefact 3.7 contains the 14 CaseMix questions that form the core of the benchmark. The benchmark ships with a compact task suite that mirrors day-to-day queries posed by clinicians, health-economists and DRG coders. Table 3 lists the nine quantitative prompts (IDs 1–9). Their answers are deterministic code sets, so an LLM can be scored automatically by exact match against the ground-truth sheet. Table 4 adds five qualitative prompts (IDs 10–14) that demand free-text justifications—e.g., drafting a technical-change request graded with a human review.
Each prompt is tagged Easy, Medium or Hard to signal the cognitive load and the number of artefacts that must be cross-referenced. Together the 14 prompts form a self-contained, reproducible test-bed for comparing prompt-engineering tricks, retrieval-augmented generation, and fine-tuning strategies on hospital-funding tasks.
Table 3: CaseMix questions with automatically verifiable answers (IDs 1–9).
Q-ID Use Case Category Difficulty CaseMix Question
1 Understanding NordDRG Definition Tables Easy Could you list all DRG groups that are related to appendectomy in all DRG versions?
2 Understanding NordDRG Definition Tables Easy Could you list all DRG groups that are related to appendectomy in the Finnish DRG version?
3 Medical Terms in Local Minority Language Easy Could you list DRG groups that might be related to Finnish word umpilisäke in the Finnish DRG version?
4 Understanding NordDRG Definition Tables Easy Could you list all DRG groups in all DRG versions where it seems that a child has been born during that particular hospital stay?
5 Medical Terms in Local Minority Language Medium Could you list DRG groups that might be related to Finnish word umpilisäke in all DRG versions from all countries? Use the drg_comb column codes and combine them into one list.
6 Surgical Procedure Logic Medium For DRG 166N which procedure codes have the connection through the procedure property? Use NCSP plus codes in your answer.
7 Connection between Definition Tables Medium Which diagnosis codes are linked to DRG 166N through diagnosis properties in all DRG versions from all countries?
8 Diagnosis Properties Medium Which diagnosis (ICD) codes are connected to diagnosis property 14X03?
9 Connection between Definition Tables Medium Which diagnosis categories and properties are listed in drg_logic for DRG group 372C?
Table 4: Qualitative CaseMix prompts that require explanatory answers (IDs 10–14).
Q-ID Use Case Category Difficulty CaseMix Question
10 Surgical Procedure Logic Medium How do you know from the drg_logic table that a patient has been in a surgery?
11 Overall Understanding and Assistance Medium Could you explain the content and purpose of all definition tables in the NordDRG system?
12 Quality Assurance Hard Read the names for the ICD codes already linked to 14X03. Then, review all diagnosis codes in the dg sheet. Which five codes are most likely missing from 14X03 based on their text descriptions?
13 Technical Changes Medium Related to case: https://nordcaseforum.easyredmine.com/issues/917. Consult the PDF How to write technical changes to NordDRG and reconsider what the screenshot’s suggestion implies.
14 Technical Changes Medium Related to case: https://nordcaseforum.easyredmine.com/issues/898. Using the same PDF, draft the required technical change so that procedure 14D51 is added to NCSP codes MADE2D, MADE2H, MADE2X, MADE32 and MADE3A, thereby grouping into DRG 914P in the ICE version; avoid unnecessary OUT rows in proc_feat.
Tables 3 and 4 are paired with the answer key in Table 5. Because all three tables share a common ID column, each “right answer” row can be matched unambiguously to its corresponding question, use-case category, and difficulty level. Together they constitute a self-contained evaluation suite: the first two tables supply the prompts, while Table 5 provides the gold-standard outputs against which LLM responses can be scored either automatically (exact-match codes) or qualitatively (open-ended explanations).
Table 5: Reference answers for each CaseMix question (IDs 1–14).
Q-ID Right answer
1 166C, 166N, 167N, 167X, 167O
2 166N, 167X, 167O
3 166N, 167X, 167O
4 370B, 370C, 370M, 370X, 371O, 371X, 372B, 372C, 372M, 372X, 373O, 373X, 374X, 375X, P05S, P05T
5 166C, 166N, 167N, 167X, 167O
6 JESA00, JESA01, JESA10, JESW96, JESW97
7 C1810, K3520
8 O6010, O6020, O6030, O6230, O6300, O6310, O6320, O6390, O6800, O6810, O6820, O6830, O6880, O6890, O7100, O7110, O7550, O7570, O8000, O8010, O8080, O8090, O8100, O8110, O8120, O8130, O8140, O8150, O8300, O8310, O8320, O8330, O8340, O8380, O8390, O8400, O8410, O8420, O8480, O8490, Z3700, Z3710, Z3720, Z3730, Z3740, Z3750, Z3760, Z3770, Z3790
9 14X03, 14X11, 14X13, 00X10
10 IN: drglogic or_proc = S; IN: proc_feat vartype = OR, varval = 1 for drg_logic procpro
11 Qualitative question
12 Qualitative question
13 Qualitative question
14 Qualitative question
4 Demonstration: Baseline LLM Performance
Table 6 benchmarks five proprietary end-points—GPT-4o, o4-mini-high, and the o3 model from OpenAI, plus Google’s Gemini 2.5 Flash and Gemini 2.5 Pro—on the nine automatically verifiable prompts (IDs 1–9 in Table 3). All systems were given the highest-fidelity bundle, Definition Tables + PDF Instructions, and an answer was marked correct only if the returned code set matched the ground-truth list exactly.
The results stratify into three tiers. The legacy o3 solved all nine tasks; GPT-4o and o4-mini-high each missed two (IDs 4 and 6, and IDs 5 and 4 respectively); Gemini 2.5 Pro and Gemini 2.5 Flash lagged further, answering 5/9 and 3/9. Performance gaps widened on cross-table reasoning tasks, underscoring why a domain-specific benchmark is needed: differences obscured by generic leaderboards become obvious once models must navigate the full NordDRG logic and its multilingual terminology.
Table 6: Accuracy of five proprietary LLM endpoints on the nine quantitative NordDRG tasks (QID 1–9) with the Definition Tables + PDF Instructions bundle.
Input Material Q-ID GPT-4o o4-mini o3 Gemini 2.5 Flash Gemini 2.5 Pro
Definition Tables + PDF Instr. 1
2
3
4
5
6
7
8
9
Table 7: Accuracy of five proprietary LLM endpoints when only the structured definition tables were provided as context. The first nine rows are the quantitative tasks (QID 1–9); the last line isolates the Finnish-only excerpt of QID 3. Compare with Table 6, where the models also received the PDF manuals.
Input Material Q-ID GPT-4o o4-mini o3 Gemini 2.5 Flash Gemini 2.5 Pro
Definition Tables 1
Definition Tables 2
Definition Tables 3
Definition Tables 4
Definition Tables 5
Definition Tables 6
Definition Tables 7
Definition Tables 8
Definition Tables 9
Definition Tables : FI subset 3
5 Evaluation
Following the Design–Science Research Methodology, we evaluate the benchmark against the three objectives stated in Section 1.5. Our strategy combines a descriptive inspection of the artefacts with an analytical baseline run.
5.1 Objective O1 — Coverage and fidelity
Dataset completeness.
[…]
Task alignment.
Nine quantitative prompts (Table 3) join several tables, while the five qualitative prompts (Table 4) reference both the definition workbook (3.1) and the expert PDFs (3.2). Thus every artefact class—tables, documentation, and prompts—is exercised at least once.
5.2 Objective O2 — Extensibility and sustainability
Schema stability and drop-in updates.
Core identifier columns—DRG, ICD, NCSP, etc.— are identical across successive Nordic releases, so the data model does not break when new rows are appended. Practically, a contributor can create a new folder in the GitHub repository, drop in the next year’s workbook plus the companion PDF manuals, and the benchmark will recognise the added version without any code changes or column renaming. This “folder-per-release” convention keeps the resource evergreen while preserving full back-compatibility for earlier experiments.
Open workflow.
All artefacts live in a GitHub repository. Adding a new task requires only a one-line entry in the prompt pack (3.7) plus an answer row in the gold file; no source code modification is needed.
5.3 Objective O3 — Baseline differentiation
The baseline in Table 6 shows a clear three-tier spread: o3 solves 9/9 tasks, GPT-4o and o4-mini-high score 7/9, while Google’s Gemini 2.5 Pro and Gemini 2.5 Flash reach 5/9 and 3/9. These gaps are largest on cross-table joins (QIDs 4–7), confirming that domain-specific logic is not captured by generic leaderboards.
Taken together, the artefact inspection (coverage, schema, contribution workflow) and the baseline run (accuracy spread) confirm that NordDRG-AI-Benchmark meets its design objectives and provides a robust yardstick for future research on LLMs in hospital funding.
6 Communication
The final DSRM step disseminates NordDRG-AI-Benchmark so that others can evaluate, critique, and extend it.
Open repository.
All artefacts, answer keys, and baseline notebooks are hosted in a public GitHub repository (https://github.com/longshoreforrest/norddrg-ai-benchmark). The repository uses a folder-per-release convention—new country versions or yearly updates can be added by copying the official workbook and PDF manuals into a dated sub-folder.
Contribution workflow.
Pull requests that add (i) new DRG versions, (ii) extra prompts, or (iii) bug-fixes to existing tables are welcomed.
Scholarly venues.
Early prototypes of the benchmark were presented at CHIRA 2024 and PCSI 2024 [Pitkäranta, 2024b, Pitkäranta, 2024a]. The current article targets a broader audience via the LNCS proceedings and a mirrored arXiv pre-print, ensuring long-term accessibility.
Together, the open repository, clear contribution guidelines, and scholarly outreach convert NordDRG-AI-Benchmark from a static data drop into a living community resource, completing the loop outlined in the DSRM methodology.
Acknowledgements
We would like to acknowledge the Nordic CaseMix Centre (NCC) for its significant contributions to the development of NordDRG-AI-Benchmark by providing access to NordDRG definition tables.
We also thank Kristiina Kahur, CEO of NCC, for her valuable comments on the manuscript, which helped improve the clarity and accuracy of this paper.